Quantitative Aptitude Practice Question and Answer

Q: A can do a work in 9 days, B can do a work in 7 days, C can do a work in 5 days. A works on the first day, B works on the second day and C on the third day respectively that is they work on alternate days. When will they finish the work ? 1651 0

  • 1
    [7 + (215/345)] days
    Correct
    Wrong
  • 2
    [6 + (11/215)] days
    Correct
    Wrong
  • 3
    [6 + (261/315)] days
    Correct
    Wrong
  • 4
    [5 + (112/351)] days
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 3. "[6 + (261/315)] days"
Explanation :

Answer: C) [6 + (261/315)] days Explanation: After day 1, A finishes 1/9 of the work. After day 2, B finishes 1/7 more of the total work. (1/9) + (1/7) is finished. After day 3, C finishes 1/5 more of total work. Total finished is 143/315. So, after day 6, total work finished is 286/315. ┬а Now remaining work =┬а 29 /315 ┬а On day 7, A will work again┬а ┬а Work will be completed on day 7 when A is working. He must finish 29/315 of total remaining work. ┬а Since he takes 9 days to finish the total task, he will need 261/315 of the day.┬а ┬а Total days required is 6 + (261/315) days.

Q: 3 : 12 :: 5 : ? 1649 0

  • 1
    17
    Correct
    Wrong
  • 2
    30
    Correct
    Wrong
  • 3
    26
    Correct
    Wrong
  • 4
    32
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 2. "30"
Explanation :

Answer: B) 30 Explanation: 332┬а+┬а3┬а=┬а12552┬а+┬а5┬а=┬а30

Q: The ratio between the length and breadth of a rectangular field is 5 : 4 if the breadth is 20m less than the length, the perimeter of the field is : 1646 0

  • 1
    260m
    Correct
    Wrong
  • 2
    280m
    Correct
    Wrong
  • 3
    360m
    Correct
    Wrong
  • 4
    none of these
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer :
Explanation :

Answer: Explanation:

Q: The difference of two numbers is 1365. On dividing the larger number by the smaller, we get 6 as quotient and 15 as remainder. What is the smaller number ┬а? 1645 0

  • 1
    270
    Correct
    Wrong
  • 2
    1270
    Correct
    Wrong
  • 3
    350
    Correct
    Wrong
  • 4
    720
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 1. "270"
Explanation :

Answer: A) 270 Explanation: Let the smallest number be x.Then larger number = (x + 1365)x + 1365 = 6x + 15= 5x = 1350 x = 270Smaller number = 270.

Q: In the given number series which follows the definite rules then find out the value of (M, N, O, and P) in second number series. 7┬а 18┬а 42┬а 94┬а 204 4┬а ┬аM┬а ┬аN┬а ┬а O┬а ┬а ┬аP What value will come in place of P? 1645 0

  • 1
    179
    Correct
    Wrong
  • 2
    184
    Correct
    Wrong
  • 3
    164
    Correct
    Wrong
  • 4
    156
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 4. "156"
Explanation :

Answer: D) 156 Explanation: Here the given series┬а7┬а 18┬а 42┬а 94┬а 204┬аfollows a pattern that,┬а 7 7┬а├Ч2+4┬а=┬а18 18┬а├Ч2+6 =┬а42 42┬а├Ч2+10 =┬а94 94┬а├Ч2+16 =┬а204 Similarly, 4 4┬а├Ч2+4┬а=┬а12 12┬а├Ч2+6 =┬а30 30┬а├Ч2+10 =┬а70 70┬а├Ч2+16 =┬а156 Hence, the value of P = 156.

Q: An automobile financier claims to be lending money at S.I., but he includes the interest every six months for calculating the principal. If he is charging an interest of 8%, the effective rate of interest becomes ? 1644 0

  • 1
    10.25 %
    Correct
    Wrong
  • 2
    8.16 %
    Correct
    Wrong
  • 3
    9.63 %
    Correct
    Wrong
  • 4
    0.16 %
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 2. "8.16 %"
Explanation :

Answer: B) 8.16 % Explanation: Let the sum be Rs. 100. Then,S.I. for first 6 months = (100 x 8 x 1) / (100 x 2) = Rs. 4S.I. for last 6 months = (104 x 8 x 1) / (100 x 2) = Rs. 4.16So, amount at the end of 1 year = (100 + 4 + 4.16) = Rs. 108.16Effective rate = (108.16 - 100) = 8.16%.

Q: If two dice are thrown together, the probability of getting an even number on one die and an odd number on the other is ┬а? 1644 0

  • 1
    1
    Correct
    Wrong
  • 2
    1/2
    Correct
    Wrong
  • 3
    0
    Correct
    Wrong
  • 4
    3/5
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 2. "1/2"
Explanation :

Answer: B) 1/2 Explanation: The number of exhaustive outcomes is 36.Let E be the event of getting an even number on one die and an odd number on the other. Let the event of getting either both even or both odd then = 18/36 = 1/2P(E) = 1 - 1/2 = 1/2.

Q: 4 men and 6 women can complete a work in 8 days, while 3 men and 7 women can complete it in 10 days. In how many days will 10 women complete it ? 1643 0

  • 1
    40 days
    Correct
    Wrong
  • 2
    36 days
    Correct
    Wrong
  • 3
    32 days
    Correct
    Wrong
  • 4
    34 days
    Correct
    Wrong
  • Show AnswerHide Answer
  • Workspace

Answer : 1. "40 days"
Explanation :

Answer: A) 40 days Explanation: Let 1 man's 1 day work = x and 1 woman's 1 day work = y.Then, 4x + 6y = 1/8 and 3x + 7y = 1/10Solving these two equations, we get:x = 11/400 and y = 1/40010 woman's 1 day work = (1/400 x 10) = 1/40. Hence, 10 women will complete the work in 40 days.

      Report Error

    Please Enter Message
    Error Reported Successfully

      Report Error

    Please Enter Message
    Error Reported Successfully

      Report Error

    Please Enter Message
    Error Reported Successfully

      Report Error

    Please Enter Message
    Error Reported Successfully

      Report Error

    Please Enter Message
    Error Reported Successfully

      Report Error

    Please Enter Message
    Error Reported Successfully

      Report Error

    Please Enter Message
    Error Reported Successfully

      Report Error

    Please Enter Message
    Error Reported Successfully